Find three consecutive even integers such that three times the middle integers is 10 more than the greatest


answer:x=2 the integers are 4,6,8
anyone tell me how my text book got this answer full break down thank you

Answer by nerdybill(5411)

Find three consecutive even integers such that three times the middle integers is 10 more than the greatest
.
Let x = smallest consecutive even integer
then
x+2 = middle of three consecutive even integers
x+4 = largest of three consecutive even integers
.
3(x+2) = (x+4) + 10
3x+6 = x+4+10
3x+6 = x+14
2x+6 = 14
2x = 8
x = 4 (smallest)
.
Middle:
x+2 = 4+2 = 6
.
Largest:
x+4 = 4+4 = 8
.
The three integers: 4,6,8